Assume you dissolve 0.235 g of the weak acid benzoic acid, C6H5CO2H, in enough water to make 5.00 ✕ 102 mL of solution and then titrate the solution with 0.128 M NaOH.

C6H5CO2H(aq) + OH-(aq) C6H5CO2-(aq) + H2O(ℓ)

What are the concentrations of the following ions at the equivalence point?

Na+, H3O+, OH-

C6H5CO2-

_____ M Na+

_____ M H3O+

_____ M OH-

_____ M C6H5CO2-

What is the pH of the solution?

pH =

please help if you can :)
Answer #1

1)

moles of benzoic acid = 0.235 / 122.12 = 1.924 x 10^-3 mol

concentration of benzoic acid = 1.924 x 10^-3 / 0.500

                                               = 3.85 x 10^-3 M

At equivalence point :

moles of acid = moles of NaOH

1.924 x 10^-3 = 0.128 x V

V = 0.015 L

total volume = 0.015 + 0.500 = 0.515 L

here salt only remains.

[C6H5COO-] = 1.924 x 10^-3 / 0.515 = 0.00374 M

C6H5COO-    +   H2O    ------------> C6H5COOH +   OH-

0.00374                                                    0                    0

0.00374 - x                                              x                        x

Kb = x^2 / 0.00374 - x

1.59 x 10^-10 = x^2 / 0.00374 - x

x = 7.705 x 10^- 7

[OH-] = 7.705 x 10^-7 M

pOH = 6.11

pH = 7.89

0.128 M   = Na+

1.30 x 10^-8 M = H3O+

7.71 x 10^-7 M   = OH-

0.00374 M = C6H5CO2-

pH = 7.89